# bayesNMF: Fast Bayesian Poisson NMF with Automatically Learned Rank Applied to Mutational Signatures

### v2

@@ 7 Discussion
In this paper, we introduced bayesNMF, a novel computationally efficient Gibbs sampler for Bayesian Poisson NMF and an automatic rank-learning approach. Through simulation studies and a large-scale data analysis, we demonstrated that bayesNMF achieves comparable accuracy and uncertainty quantification to standard Bayesian Poisson NMF at a fraction of the computational cost. All methods are available in an open-source R software package on GitHub at jennalandy/bayesNMF.
The Metropolis-Hastings steps are the key to bayesNMF’s computational efficiency. In future work, these updates could easily be used in conjunction with other approaches, including compressive NMF [Zito and Miller, 2024] or multistudy NMF [Grabski et al., 2023b]. This concept of model-paired proposals may also be generalizable to other hierarchical count models that rely on Poisson augmentation, for example Poisson regression with random effects.
We recognize a potential limitation of our assumption that each estimated signature must be assigned to a unique reference signature. In some cases, precision reported in Figure 3 could be higher if multiple estimated signatures were allowed to map to the same reference. Similarly, some SignatureAnalyzer-estimated signatures in the PCAWG analysis that show cosine similarity to their assigned references (Appendix Figure D.2) may have achieved higher similarity if duplicate assignments were permitted. This is only a limitation of our downstream analysis and visualization tools, not of the models themselves.
In summary, bayesNMF offers a powerful toolbox for exploring complex mutational landscapes, especially with its ability to report posterior uncertainty while learning rank as part of the Bayesian model. The open source R software package encourages usability of these methods and reproducibility of our results. As cancer genomics databases expand, bayesNMF’s efficiency and flexibility make it well suited to drive further research.
@@ 8 Acknowledgments
Appendices:
A) Model specifications (likelihoods, priors, hyperpriors) and Gibbs updates. B) Implementation details (convergence, warmup, inference, reference alignment, label switching diagnostic). C) Additional simulation study results, including sensitivity analyses for sparse data and changing rank range. D) Additional PCWAG data application results. (AppendixA.pdf, etc.)
@@
The R package bayesNMF is publicly available on GitHub at jennalandy/bayesNMF. Documented by README.md and vignettes.
@@
All code needed to reproduce simulation studies and data analysis is publicly available on GitHub at jennalandy/bayesNMF_PAPER, documented by README.md.
@@
The PCAWG data is available through the ICGC ARGO data portal: access instructions link. Instructions are also included in the code file 1_pcawg_data.qmd.
@@ Appendix A Appendix A: Model Specifications
Likelihood Prior Sampler # Updates # Acceptance Ratios Section
@@
Table A.3: List of models implemented in bayesNMF, reporting the number of parameter/prior parameter updates and the number of acceptance rates computed per iteration of the Gibbs sampler. Both values directly impact compute time and memory consumption. Typically $N<<K,G$. MH-within-Gibbs samplers are the new contribution of this work.
@@ A.1.1 Model Definition
Priors: $P_{kn}\sim\text{TruncNormal}_{[0,\infty)}\left(\mu^{P}_{kn},(\sigma^{P}_{kn})^{2}\right)$
$E_{ng}\sim\text{TruncNormal}_{[0,\infty)}\left(\mu^{E}_{ng},(\sigma^{E}_{ng})^{2}\right)$
@@
The prior and hyperprior structure on $A$ is defined in Section A.0.1. From here on, let $\theta$ to refer to all prior and hyperprior parameters, and assume any use of “TruncNorm” uses the bounds $0,\infty)$.
@@ A.1.4 Parameter Gibbs Updates: MH-Within-Gibbs steps
$f(P^{*}_{kn}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}\left(P^{*}_{kn}|\mu^{P}_{kn},(\sigma^{P}_{kn})^{2}\right)\prod_{g}p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)$
$f(E^{*}_{ng}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}\left(E^{*}_{ng}|\mu^{E}_{ng},(\sigma^{E}_{ng})^{2}\right)\prod_{k}p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE^{*})_{kg}\right)$
@@
$g(P^{*}_{kn}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}(P^{*}_{kn}|\mu^{P}_{kn},(\sigma^{P}_{kn})^{2})\prod_{g}p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)$
$\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{P}_{kn}(M,P,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,P,E,\theta)\right)$
@@
$g(E^{*}_{ng}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}(E^{*}_{ng}|\mu^{E}_{ng},(\sigma^{E}_{ng})^{2})\prod_{k}p_{\text{Normal}}\left(M_{kg}|\mu=(PE^{*})_{kg},\sigma^{2}=(PE)_{kg}\right)$
$\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{E}_{ng}(M,P,A,E,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,E,\theta)\right)$
@@ A.2.4 Parameter Gibbs Updates: MH-Within-Gibbs steps
$\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{P}_{kn}(M,P,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,P,E,\theta)\right)$
@@
$\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{E}_{ng}(M,P,A,E,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,E,\theta)\right)$
@@ A.5.1 Model Definition
Priors: $P_{kn}\sim\text{TruncNormal}_{[0,\infty)}\left(\mu^{P}_{kn},(\sigma^{P}_{kn})^{2}\right)$
$E_{ng}\sim\text{TruncNormal}_{[0,\infty)}\left(\mu^{E}_{ng},(\sigma^{E}_{ng})^{2}\right)$
@@
The prior and hyperprior structure on $A$ is defined in Section A.0.1. From here on, let $\theta$ to refer to all prior and hyperprior parameters, and assume any use of “TruncNorm” uses the bounds $0,\infty)$.
@@ A.5.3 Parameter Gibbs Updates
$P_{kn}|$... $\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{P}_{kn}(M,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,A,E,\theta)\right)$
@@
$E_{ng}|$... $\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{E}_{ng}(M,P,A,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,\theta)\right)$
@@ A.6.3 Parameter Gibbs Updates
$P_{kn}$... $\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{P}_{kn}(M,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,A,E,\theta)\right)$
@@
$E_{ng}|$... $\sim\text{TruncNorm}_{[0,\infty)}\left(\mu=m^{E}_{ng}(M,P,A,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,\theta)\right)$
@@ B.4 Alignment to Reference Signatures
We follow existing mutational signatures works [Rosales et al., 2016, Gori and Baez-Ortega, 2018, Alexandrov et al., 2020, Grabski et al., 2023b] and focus on the scale-independent measure of cosine similarities between signatures. To align estimated factors to the reference factors, we first create a similarity matrix $S$, where each element is the cosine similarity between a column of $P$ and a column of $\hat{P}$, $S_{ij}=cos(P_{i},\hat{P}_{j})$. The Hungarian algorithm [Kuhn, 1955] is used on $1\cdot S$, which solves the combinatorial optimization problem to reorder rows and columns to minimize the sum of values on the diagonal. The resulting aligned similarity matrix, $S^{*}$, has an optimal alignment in terms of total aligned cosine similarity. This approach only allows a single estimated signature to be aligned to each reference signature, and vice versa.
@@ D.2 Additional Results
For PCAWG results using bayesNMF+SBFI and SignatureAnalyzer+ARD, Figure D.2 reports median number of mutations attributed to each signature and the cosine similarity of each estimated signature to its assigned reference signature (posterior average for bayesNMF).
Table D.5 reports estimated rank for PCAWG results using bayesNMF+SBFI and SignatureAnalyzer+ARD.
Figure D.3 investigates signatures assigned by both bayesNMF and SignatureAnalyzer versus signatures assigned by bayesNMF only. The signatures assigned by bayesNMF only have lower cosine similarity to the assigned reference and a lower proportion of assignment votes, indicating uncertainty in assignment.

### v3

@@ 7 Discussion
We introduced bayesNMF, a MH-within-Gibbs sampler for Bayesian Poisson NMF with automatic rank-learning. The method is built on two primary innovations: an efficient MH-within-Gibbs sampling strategy to avoid Poisson augmentation, and a BIC-based sparsity prior that enables rank to be learned accurately within a single Bayesian formulation. Simulation studies and a large-scale data analysis demonstrate that bayesNMF achieves comparable accuracy and uncertainty quantification to standard Bayesian Poisson NMF at a fraction of the computational cost. All methods are available in an open-source R software package on GitHub at jennalandy/bayesNMF.
The MH steps with our high-overlap proposals are the key to bayesNMF’s computational efficiency. In future work, these updates could be used in conjunction with other approaches such as compressive NMF [Zito and Miller, 2024] or multistudy NMF [Grabski et al., 2025]. Additionally, our model-paired proposals may be generalizable to other hierarchical count models that rely on Poisson augmentation, like Poisson regression with random effects.
We recognize a potential limitation of our assumption that each estimated signature must be assigned to a unique reference signature. In some cases, the precision reported in Figure 3 could be higher if multiple estimated signatures were allowed to map to the same reference. Similarly, some SignatureAnalyzer-estimated signatures in the PCAWG analysis (Appendix Figure D.2) may have achieved higher similarity to their assigned references if duplicate assignments were permitted. However, this is only a limitation of our downstream analysis and visualization tools, not of the models themselves.
In summary, bayesNMF offers a powerful toolbox for exploring complex mutational landscapes, especially with its ability to report posterior uncertainty while learning rank as part of the Bayesian model. The open source R software package facilitates the usability of these methods and the reproducibility of our results. As cancer genomic databases expand, bayesNMF’s efficiency and flexibility make it well-suited to drive further research.
@@ 8 Acknowledgments
Appendix:
A) Model specifications (likelihoods, priors, hyperpriors) and Gibbs updates. B) Implementation details (convergence, warmup, inference, reference alignment, label switching diagnostic). C) Additional simulation study results, including sensitivity analyses for sparse data and increasing rank range. D) Additional PCAWG data application results.
@@
The R package bayesNMF is publicly available on GitHub at jennalandy/bayesNMF, documented by the README and vignettes.
@@
All code needed to reproduce simulation studies and data analysis is publicly available on GitHub at
jennalandy/bayesNMF_PAPER, documented by README.md.
@@
The PCAWG data are available through the ICGC ARGO data portal:
access instructions at docs.icgc-argo.org/docs/data-access/. Instructions are also included in the code file 1_pcawg_data.qmd.
@@ Appendix A Appendix A: Model Specifications
We introduce all models using the simpler prior on the signature inclusion matrix $A$ corresponding to Bayesian Factor Inclusion (BFI). See the main text for details on the prior of $A$ under Sparse BFI (SBFI). If rank is fixed, $A$ is set as an identity matrix and Gibbs updates of $A$ are skipped.
@@
Likelihood Prior Sampler # Updates # Acc. Ratios Sec.
@@
Table A.3: List of models implemented in bayesNMF, reporting the number of parameter/prior parameter updates and the number of acceptance rates computed per iteration of the Gibbs sampler. Both values directly impact compute time and memory consumption. Typically $N<<K,G$.
@@ A.1.1 Model Definition
Priors: $P_{kn}\sim\text{TruncNorm}\left(\mu^{P}_{kn},(\sigma^{P}_{kn})^{2},0,\infty\right)$
$E_{ng}\sim\text{TruncNorm}\left(\mu^{E}_{ng},(\sigma^{E}_{ng})^{2},0,\infty\right)$
@@
The prior and hyperprior structure on $A$ is defined in Section A.0.1. From here on, let $\theta$ to refer to all prior and hyperprior parameters, and assume the parameter order for Truncated Normal distributions $\text{TruncNorm}(\mu,\sigma^{2},a,b)$ where $a$ and $b$ are truncation bounds.
@@ A.1.4 Parameter Gibbs Updates: MH-Within-Gibbs steps
$f(P^{*}_{kn}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}\left(P^{*}_{kn}|\mu^{P}_{kn},(\sigma^{P}_{kn})^{2},0,\infty\right)\prod_{g}p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)$
$f(E^{*}_{ng}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}\left(E^{*}_{ng}|\mu^{E}_{ng},(\sigma^{E}_{ng})^{2},0,\infty\right)\prod_{k}p_{\text{Poisson}}\left(M_{kg}|\lambda=(PE^{*})_{kg}\right)$
@@
$g(P^{*}_{kn}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}(P^{*}_{kn}|\mu^{P}_{kn},(\sigma^{P}_{kn})^{2},0,\infty)\prod_{g}p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)$
$P^{*}_{kn}|P,E,M,\theta$ $\sim\text{TruncNorm}\left(\mu=m^{P}_{kn}(M,P,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,P,E,\theta),0,\infty\right)$
@@
$g(E^{*}_{ng}|P,E,M,\theta)$ $\propto p_{\text{TruncNorm}}(E^{*}_{ng}|\mu^{E}_{ng},(\sigma^{E}_{ng})^{2},0,\infty)\prod_{k}p_{\text{Normal}}\left(M_{kg}|\mu=(PE^{*})_{kg},\sigma^{2}=(PE)_{kg}\right)$
$E^{*}_{ng}|P,E,M,\theta$ $\sim\text{TruncNorm}\left(\mu=m^{E}_{ng}(M,P,A,E,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,E,\theta),0,\infty\right)$
@@ A.2.4 Parameter Gibbs Updates: MH-Within-Gibbs steps
$P^{*}_{kn}|P,E,M,\theta$ $\sim\text{TruncNorm}\left(\mu=m^{P}_{kn}(M,P,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,P,E,\theta),0,\infty\right)$
@@
$E^{*}_{ng}|P,E,M,\theta$ $\sim\text{TruncNorm}\left(\mu=m^{E}_{ng}(M,P,A,E,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,E,\theta),0,\infty\right)$
@@ A.5.1 Model Definition
Priors: $P_{kn}\sim\text{TruncNorm}\left(\mu^{P}_{kn},(\sigma^{P}_{kn})^{2},0,\infty\right)$
$E_{ng}\sim\text{TruncNorm}\left(\mu^{E}_{ng},(\sigma^{E}_{ng})^{2},0,\infty\right)$
@@
The prior and hyperprior structure on $A$ is defined in Section A.0.1. From here on, let $\theta$ to refer to all prior and hyperprior parameters, and assume the parameter order for Truncated Normal distributions $\text{TruncNorm}(\mu,\sigma^{2},a,b)$ where $a$ and $b$ are truncation bounds.
@@ A.5.3 Parameter Gibbs Updates
$P_{kn}|$... $\sim\text{TruncNorm}\left(\mu=m^{P}_{kn}(M,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,A,E,\theta),0,\infty\right)$
@@
$E_{ng}|$... $\sim\text{TruncNorm}\left(\mu=m^{E}_{ng}(M,P,A,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,\theta),0,\infty\right)$
@@ A.6.3 Parameter Gibbs Updates
$P_{kn}$... $\sim\text{TruncNorm}\left(\mu=m^{P}_{kn}(M,A,E,\theta),\sigma^{2}=s^{P}_{kn}(M,A,E,\theta),0,\infty\right)$
@@
$E_{ng}|$... $\sim\text{TruncNorm}\left(\mu=m^{E}_{ng}(M,P,A,\theta),\sigma^{2}=s^{E}_{ng}(M,P,A,\theta),0,\infty\right)$
@@ B.4 Alignment to Reference Signatures
We follow existing mutational signatures works [Rosales et al., 2016, Gori and Baez-Ortega, 2018, Alexandrov et al., 2020, Grabski et al., 2025] and focus on the scale-independent measure of cosine similarities between signatures. To align estimated factors to the reference factors, we first create a similarity matrix $S$, where each element is the cosine similarity between a column of $P$ and a column of $\hat{P}$, $S_{ij}=cos(P_{i},\hat{P}_{j})$. The Hungarian algorithm [Kuhn, 1955] is used on $1\cdot S$, which solves the combinatorial optimization problem to reorder rows and columns to minimize the sum of values on the diagonal. The resulting aligned similarity matrix, $S^{*}$, has an optimal alignment in terms of total aligned cosine similarity. This approach only allows a single estimated signature to be aligned to each reference signature, and vice versa.
@@ D.2 Additional Results
  • Figure D.2: For PCAWG results using bayesNMF+SBFI and SignatureAnalyzer+ARD, this figure reports median number of mutations attributed to each signature and the cosine similarity of each estimated signature to its assigned reference signature (posterior average for bayesNMF).
  • Table D.5: Reports estimated rank for PCAWG results using bayesNMF+SBFI and SignatureAnalyzer+ARD.
  • Figure D.3: Investigates signatures assigned by both bayesNMF and SignatureAnalyzer versus signatures assigned by bayesNMF only. The signatures assigned by bayesNMF only have lower cosine similarity to the assigned reference and a lower proportion of assignment votes, indicating uncertainty in assignment.